OpenStudy (anonymous):

solve the equation 111 5 ----- = ---- 20 4

OpenStudy (blurbendy):

is there a variable you have to solve for?

OpenStudy (anonymous):

no

OpenStudy (blurbendy):

eh, 111/20 = 5/4 is a false statement

OpenStudy (anonymous):

well the 111 part, is just 3 lines that look like ones, i dont know if thats a variable or?

OpenStudy (blurbendy):

i thought it was the number one hundred and eleven

OpenStudy (anonymous):

omg its an m i think im blind, hahha sorry

OpenStudy (johnweldon1993):

haha...so it's \[\large \frac{ m }{ 20 } = \frac{ 5 }{ 4 }\] right?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

So you would cross multiply....and then solve for 'm' what do you get after you cross multiply?

OpenStudy (anonymous):

so it would be 5 * 20 divided by 4?

OpenStudy (johnweldon1993):

That would be correct...which comes out to...

OpenStudy (johnweldon1993):

Great job!...that would be 'm'

OpenStudy (anonymous):

thank you(:

OpenStudy (johnweldon1993):

No problem! :)